OpenStudy (anonymous):

What is the solution to the equation log4 (x) – log4 (2) = 3 ?

OpenStudy (anonymous):

log4(x/2)=3 x/2=4^3 x=32

OpenStudy (anonymous):

log4(x/2)=3 x/2=4^3 x=128